OpenStudy (anonymous):

Nanna's Italian Restaurant has collected data about customer sauce orders. It calculated that P(marinara) = 0.64, P(alfredo) = 0.56, and P(marinara or alfredo) = 0.89. Determine the P(marinara and alfredo).

OpenStudy (anonymous):

I'm not understanding this, if .89 is both than i figure the marinara and alfredo would be 11% or .11 but that's not one of the answers

OpenStudy (anonymous):

P(m and a) = P(m) + P(a) - P(m or a) Use that relationship to solve your problem.

OpenStudy (anonymous):

That's a general relationship that is always true. It doesn't matter what A and B are Pr(A and B) = Pr(A) + Pr(B) - Pr(A or B) and Pr(A or B) = Pr(A) + Pr(B) - Pr(A and B)

OpenStudy (anonymous):

Uhm, that doesn't make sense. Could you explain? If P (m+a)= .89 then doing P(m or A) lets say A would just leave me with the original m

OpenStudy (anonymous):

Or is that not how it works?

OpenStudy (ranga):



OpenStudy (anonymous):

I was just about to draw something like the above.

OpenStudy (anonymous):

OOOH okay so taking away the B from B & A would only take away the B included in A?

OpenStudy (anonymous):

Yes, it makes sense when you look at the diagram. Just tuck this formula away for future use , though, so you don't have to derive it every time\ Pr(A and B) = Pr(A) + Pr(B) - Pr(A or B) and Pr(A or B) = Pr(A) + Pr(B) - Pr(A and B)\ It is VERY useful

OpenStudy (anonymous):

So what answer do you get for your problem?

OpenStudy (anonymous):

I got .31 or 31%

OpenStudy (anonymous):

Is this correct? o:

OpenStudy (anonymous):

yes. that's what i got, too

OpenStudy (anonymous):

cool! Thank you! ^_^

OpenStudy (anonymous):

No problem